Hi - we are having an issue turning our 5977B MS. The problem started about a month ago (tune file attached). We changed to filament 2 and the problem went away for a few weeks. It is back again now. Since the filament appeared to be the solution last time, we replaced both filaments yesterday (4/3) . Neither of the new filaments are providing acceptable tunes. To make this even more confusing, the tune reports themselves are inconsistent. I've also attached a tune report that appears to be somewhat acceptable (although the 502 mass response is low). Now when we tune the MS the generated reports look like either of the attachments, without changing anything (simply running one tune immediately after the next).
Any idea why the instrument is not tuning consistently? Ideas for next steps?